# Symmetrizing the Constraints – Density Matrix Renormalization Group for Constrained Lattice Models

### v1

@@ Abstract
We develop a density matrix renormalization group (DMRG) algorithm for constrained quantum lattice models that successfully implements the local constraints as symmetries in the contraction of the matrix product states and matrix product operators. Such an implementation allows us to investigate a quantum dimer model in DMRG with substantial circumference on cylindrical geometry for the first time in the literature. We have thence computed the ground state phase diagram of the quantum dimer model on triangular lattice, with the symmetry-breaking characteristics of the columnar solid phase and $\sqrt{12}\times\sqrt{12}$ valence bond solid phase fully captured, as well as the topological entanglement entropy of the $\mathbb{Z}_{2}$ quantum spin liquid phase that extends to the RK point on non-bipartite lattice accurately revealed. Our DMRG algorithm on constrained quantum lattice models opens new opportunities for matrix and tensor-based algorithms for these systems that have immediate relevance towards the frustrated quantum magnets and synthetic quantum simulators.
@@ I Introduction
Based on the description above, it can be seen that previous studies employing ED and QMC methods have demonstrated significant difficulties due to the frustrations and constraints in the QDM model. ED works only for small clusters. QMC simulations, on the other hand, can be quite challenging, especially when addressing the highly degenerate states that encompass all maximally flippable plaquette configurations. Previous works based on the zero-temperature Green’s function QMC [21, 22, 27, 23] and path-integral-based sweeping cluster QMC [24] on triangular lattice QDM, require extensive sampling and either very long projection or very low temperatures to obtain meaningful results in columnar, $\sqrt{12}\times\sqrt{12}$ VBS and the $\mathbb{Z}_{2}$ QSL phases. The computational cost scales rapidly with system size and the criticality of the VBS-QSL transition is still not fully determined (see below).
@@
It is under such circumstances, we find it interesting that the matrix product state (MPS) or tensor-based methods have not been fully applied to constrained lattice models. To the best of our knowledge, the only density matrix renormalization group (DMRG) attempts for constrained model are done either by selecting the good basis states that satisfy the constraint at each level of the algorithm and filtering out the states that do not satisfy the QDM constraints [56, 57, 58], or by adding additional potential terms to penalize states that violates the dimer constraints [59]. Both practices, of course, work when the system is small and 1D-like (e.g., a ladder or cylinder with small circumference), but will become cumbersome when approaching the 2D limit. The important property of MPS and matrix product operator (MPO) in DMRG, that is, the symmetry of the model, has not been exploited in previous attempts for the QDM, certainly not to the level of the symmetry implementation in the MPS and MPO for spin and fermion models with only global U(1) or SU(2) symmetries. The conversion of the local constraint into symmetry implementation is where our development is built upon.
@@
In this work, we introduce a new DMRG algorithm designed to study systems with local constraints, and we successfully apply it to the triangular QDM on cylinder geometries. In contrast to the previous approach in dealing with states violating the constraint afterward, our approach symmetrizes the constraints, in that it treats the constraint as a set of local symmetries and implements them directly within the DMRG framework to block diagonalize all local tensors. This method inherently satisfies the constraints and significantly accelerates the computation, making it applicable to any 2D system with local constraints, with the same computational complexity of DMRG for 2D systems with global symmetries. By combining this method with ED analysis, we study the ground-state phase diagram of the QDM on the triangular lattice. Our results with wide cylinders ($L_{y}=4,6,8$) successfully capture the columnar solid phase, $\sqrt{12}\times\sqrt{12}$ valence bond solid phase and the $\mathbb{Z}_{2}$ quantum spin liquid phase, at different parameter regions of the ground state phase diagram and are fully consistent with previous works on the problem with other methods [9, 20, 10, 11, 24, 21, 22, 23]. Furthermore, we investigate the EE in both the $\mathbb{Z}_{2}$ topological QSL phase and the degenerate columnar phase, and for the first time, reveal the topological EE of $\gamma=-\ln(2)$ inside the QSL phase at parameters deviated from the RK point, and the finite ground state degeneracy of the columnar phase. These results are beyond the existing literature and point out the great potential of our DMRG algorithm in studying other constrained quantum many-body systems such as quantum loop models and synthetic quantum simulators.
@@
The rest of the paper is organized as follows. In Sec. II, the main idea of our method is given, focusing on the implementation of the symmetrization of local constraints and dynamic storage of quantum numbers. Sec. III discusses our results on triangular lattice QDM, where Sec. III.1 shows the benchmark with ED for exact results and Sec. III.2 gives the structure factor of the dimer correlation function based on wide cylinders that exhibit the columnar phase, the $\sqrt{12}\times\sqrt{12}$ phase and the $\mathbb{Z}_{2}$ QSL phase in the ground state phase diagram of the model, and Sec. III.3, on the other hand, focuses on the EE results from our DMRG and for the first time in literature, we show the topological EE inside the $\mathbb{Z}_{2}$ QSL away from the exactly solvable RK point, we also show that our EE computation successfully reveals the ground state degeneracy inside the columnar phase. Finally, Sec. IV reiterates the reason for the success in our symmetrization of the local constraints and points out other interesting constrained quantum lattice models that are either of fundamental theoretical importance or of immediate experimental relevance in various quantum simulators that are awaiting to be explored with our algorithm.
@@
Figure 1: QDM on a plaquette as an algorithmic demonstration. (a) The system with black (red) numbers indexing the bonds (sites). (b) The Hamiltonian in terms of local tensors. (c) Schematic of the quantum numbers flows for the MPO in (b). The four numbers are the quantum numbers (number of dimers on each site), and each four numbers set label a virtual bond space. (d) A $6\times 6$ triangular lattice with OBC (PBC) along $x$- ($y$-) direction. The local degree of freedom lives on every bond, and they are labeled according to the DMRG path. For a two-site update on bonds 51 & 52 (blue), only 8 quantum numbers (sites in red) are active.
@@ II Method
where the $t$-term flips parallel dimers to other directions in the plaquette and the positive (negative) $V$-term repulse (attracts) parallel dimers. The RK point located at $t=V$ is an exact-solvable point, where the ground state wavefunction consists of the equal superposition of all possible dimer configurations under the constraint [37, 8].
Note that every term in the Hamiltonian commutes with the operator counting the number of dimers connected to site $v$. Although each kinetic ($t$) term swaps dimer on bonds with different orientations, the four sites it touches are connected to the same numbers of dimers before and after. Consequently, the number of dimer connected to each site is a conserved quantity, i.e., the local constraint, and we have $N_{\textrm{site}}$ many of them defined on the lattice. One can then implement these conserved quantities in the DMRG computation, the same as how one considers the total $S^{z}$ in the Heisenberg model as a good quantum number. Once these symmetries are implemented, one can speed up all computation by making use of the sparse block structure of local tensors in MPO and MPS, and the constraint will be set by the initial MPS and unchanged throughout the computation.
Moreover, the Hamiltonian commutes with operator $M=\prod_{i\in l_{M}}(-1)^{n_{i}}$ leading to two topological sectors, where $l_{M}$ is one horizontal path cutting through the whole system (red dashed line in Fig. 3 (b) & (c)). The commutation relation follows from the fact that every plaquette operator in the Hamiltonian touches either 0 or 2 bonds on $l_{M}$, leaving $M$ unchanged. Since we only have PBC along the $y$-direction, this geometry holds only two topological sectors characterized by $M=\pm 1$. This global symmetry can also be implemented in DMRG by tracking a quantity that changes sign upon crossing a dimer located on bonds along $l_{M}$.
@@
$H=J\sum_{\langle i,j\rangle}\mathbf{S}_{i}\cdot\mathbf{S}_{j}=J\sum_{\langle i,j\rangle}\{S^{z}_{i}S^{z}_{j}+\frac{1}{2}(S^{+}_{i}S^{-}_{j}+S^{-}_{i}S^{+}_{j})\}$ (2)
@@
In an MPO, the charge flow is marked when constructing the Hamiltonian MPO in the automata process [60, 61, 62]. Suppose one would like to put in the interaction $S^{+}_{2}S^{-}_{3}$, where 2 and 3 are two sites of a Heisenberg plaquette in Eq. (2). In the second local tensor, an $S^{+}_{2}$ is put in a block labeled 0 and +1 for charges flowing in and out in the virtual bond spaces, respectively, and $S^{-}_{3}$ is put in the third local tensor, with labels +1 and 0. Since all the terms in the Hamiltonian preserve such symmetry, every term will end with charge 0. By labeling flows for all dimensions like this, all local tensors are block diagonalized, where nonzero elements only appear in blocks where charges flowing in and out are the same, thus, SVD can be performed on each block successively. For the MPS, the charge flow in each local tensor is marked similarly, and the charge flowing out from the last tensor is the total $S^{z}$ for the state.
@@
For QDM or QLM, as mentioned before, the Hamiltonian preserves the number of dimers connected to each site. Therefore, we can label the MPO and MPS similarly to how it is done in the Heisenberg model, with each virtual bond space being labeled with $N_{\mathrm{site}}$ quantum numbers. In the end, for the last tensor of MPO, it will flow to $(0,0,\cdots,0)$, similar to the Heisenberg case. For the MPSs, the last flowing out should represent the number of dimers connected to each site in the state, which should be $(1,1,\cdots,1)$ for QDM and $(2,2,\cdots,2)$ for QLM, respectively.
@@
where Pauli matrices $\sigma_{i}^{+}$, $\sigma^{-}_{i}$ acting on the two-dimensional local Hilbert space on the $i$-th bond creates and annihilates a dimer respectively, and $\hat{n}_{i}$ counts the number of dimers on the $i$-th site. The MPO train formulation for this Hamiltonian is shown in Fig. 1 (b). One can recover the Hamiltonian by contracting all virtual bond dimensions.
@@
Consider the first term in the Hamiltonian $-t\sigma_{1}^{+}\sigma_{2}^{+}\sigma_{3}^{-}\sigma_{4}^{-}$, which is carried in the second channels (columns) in local tensors $M_{1}$ to $M_{3}$. $-t\sigma_{1}^{+}$ acting on the first bond creates a dimer here, thus increase the quantum numbers (numbers of dimer) on sites 1 and 2 (Fig. 1 (a)) by one, thus leads to quantum numbers flowing out at the second channel being (1,1,0,0). In $M_{2}$, $\sigma_{2}^{+}$ is placed at the (2,2) block, which means the quantum numbers flowing in is (1,1,0,0) (the quantum numbers flowing out of the second channel in $M_{1}$). A creation on bond 2 will increase quantum numbers on sites 3 and 4 by 1, thus leading to quantum numbers flowing out at the second channel being (1,1,1,1). One can follow any of the terms and check that all quantum numbers indeed flow to (0,0,0,0) eventually (Fig. 1 (c)), as guaranteed by the symmetries of Hamiltonian.
In a larger system with more sites or a more complicated lattice structure, the same symmetry can be implemented. However, in practice, one does not need to keep track of all $N_{\textrm{site}}$ many quantum numbers for the whole process, but only the active ones, which are the ones touched by previous bonds at least once and will be touched later. Consider a $6\times 6$ triangular lattice, as shown in Fig. 1 (d), if one is performing a two-site update on local tensors 51 & 52, only 8 quantum numbers (sites in red) are active. All the others are either finished before bond 51 and cannot be modified anymore or are yet to be touched until after bond 52, and the corresponding quantum numbers are all 0. For DMRG simulation on a 2D lattice wrapped around a cylinder, the number of such non-trivial quantum numbers to be stored at each local tensor can be reduced to roughly the number of sites along the periodic boundary condition ($L_{y}$).
@@ III Results
It is expected that at negative $V$, the dimers will form a columnar phase with translational symmetry breaking in one of the three lattice directions (related by 120 rotation) and 2-fold degeneracy in each of the symmetry-breaking directions. The formation of the columnar phase is due to the interplay of the attractive $V$ and an order-by-disorder effect [63, 9, 11, 64] (to overcome the extensive degeneracy $\exp(L_{y})$ generated by sliding each row of dimers at the ordered direction freely). As $V$ moves towards positive values, the columnar phase will transit into a valence bond solid (VBS) with $\sqrt{12}\times\sqrt{12}$ unit cell [11]. Such an enlarged unit cell contains 3 resonating rhombi, i.e., each unit cell contains 12 lattice sites, and it is supposed to have a Bragg peak of the dimer-dimer correlation at the $X$ point of the Brillouin zone (BZ). The transition between columnar and VBS phase is suggested to be first order [11, 21].
@@
Figure 4: Dimer correlation of QDM on triangular lattice. (a) Dimer correlation function $S(\mathbf{k})$ along the high symmetry path (denoted as the black path in panels (b-d)) for $V=-1$, $V=-0.2$ and $V=0.9$. The system size is $L_{x}=24$ and $L_{y}=6$. (b-d) are the correlation in the entire BZ. Inside the columnar phase with $V=-1$, the Bragg peak of the order is at the $M$ points (red circle); inside the $\sqrt{12}\times\sqrt{12}$ VBS phase with $V=-0.2$, the order wave vector is at $X$ points (red corss), and the $M$ point becomes secondary; inside the $\mathbb{Z}_{2}$ QSL phase with $V=0.9$, the dimer correlation is flat as the dimer (as well as vison) correlation are all short-ranged in real-space [24]. Panels (b-d) share the same colorbar on the right.

### v3

@@ Abstract
We develop a density matrix renormalization group (DMRG) algorithm for constrained quantum lattice models that successfully implements the local constraints as symmetries in the contraction of the matrix product states and matrix product operators. Such an implementation allows us to investigate a quantum dimer model in DMRG for any lattice geometry wrapped around a cylinder with substantial circumference. We have thence computed the ground state phase diagram of the quantum dimer model on triangular lattice, with the symmetry-breaking characteristics of the columnar solid phase and $\sqrt{12}\times\sqrt{12}$ valence bond solid phase fully captured, as well as the topological entanglement entropy of the $\mathbb{Z}_{2}$ quantum spin liquid phase that extends to the RK point on non-bipartite lattice accurately revealed. Our DMRG algorithm on constrained quantum lattice models opens new opportunities for matrix and tensor-based algorithms for these systems that have immediate relevance towards the frustrated quantum magnets and synthetic quantum simulators.
@@ I Introduction
Based on the description above, it can be seen that previous studies employing ED and QMC methods have demonstrated significant difficulties due to the frustrations and constraints in the QDM model. ED works only for small clusters. QMC simulations, on the other hand, can be quite challenging, especially when addressing the highly degenerate states that encompass all maximally flippable plaquette configurations. Previous works based on the zero-temperature Green’s function QMC [21, 22, 27, 23] and path-integral-based sweeping cluster QMC [24] on triangular lattice QDM, require extensive sampling and either very long projection or very low temperatures to obtain meaningful results in columnar, $\sqrt{12}\times\sqrt{12}$ VBS and the $\mathbb{Z}_{2}$ QSL phases. The computational cost scales rapidly with system size, and the criticality of the VBS-QSL transition is still not fully determined (see below).
@@
It is under such circumstances, we find it interesting that the matrix product state (MPS) or tensor-based methods have not been fully applied to constrained lattice models. To the best of our knowledge, previous density matrix renormalization group (DMRG) attempts for constrained models have been done by projecting to the good basis states that satisfy the constraint at each level of the algorithm and filtering out the states that do not satisfy the QDM constraints [56, 57, 58]; by adding additional potential terms to penalize states that violates the dimer constraints [59]; or by introducing redundancy in local physical space to take care of the constraint and having local gauge symmetry to compensate the redundancy [60]. All practices work well when the system is 1D-like (e.g., a ladder or cylinder with small circumference) or has a simpler lattice structure (square lattices), but will become cumbersome when approaching the 2D limit and for models with constraints that are hard to cure by local redundancy. The last approach introduced the idea of including local symmetry to avoid projection or penalty terms, following the common practice in spin and fermion models with global symmetry such as $\mathrm{U}(1)$ or $\mathrm{SU}(2)$ symmetries [61, 62, 63, 64, 65, 66]. The direct conversion of the local constraint into symmetry implementation is where our development is built upon.
@@
In this work, we introduce a new DMRG design to study systems with local constraints, and we successfully apply it to the triangular QDM on cylinder geometries. Our approach symmetrizes the constraints, in that it treats the constraint as a set of local symmetries and implements them directly within the DMRG framework to block diagonalize all local tensors. This method inherently satisfies the constraints and significantly accelerates the computation, making it applicable to any 2D system with local constraints, with the same computational complexity of DMRG for 2D systems with global symmetries. By combining this method with ED analysis, we study the ground-state phase diagram of the QDM on the triangular lattice. Our results with wide cylinders ($L_{y}=4,6,8$) successfully capture the columnar solid phase, $\sqrt{12}\times\sqrt{12}$ valence bond solid phase and the $\mathbb{Z}_{2}$ quantum spin liquid phase, at different parameter regions of the ground state phase diagram and are fully consistent with previous works on the problem with other methods [9, 20, 10, 11, 24, 21, 22, 23]. Furthermore, we investigate the EE in both the $\mathbb{Z}_{2}$ topological QSL phase, and for the first time, reveal the topological EE of $\gamma=-\ln(2)$ inside the QSL phase at parameters deviated from the RK point. These results are beyond the existing literature and point out the great potential of our DMRG algorithm in studying other constrained quantum many-body systems such as quantum loop models and synthetic quantum simulators.
@@
The rest of the paper is organized as follows. In Sec. II, the main idea of our method is given, focusing on the implementation of the symmetrization of local constraints and dynamic storage of quantum numbers. Sec. III discusses our results on triangular lattice QDM, where Sec. III.1 shows the benchmark with ED for exact results and Sec. III.2 gives the structure factor of the dimer correlation function based on wide cylinders that exhibit the columnar phase, the $\sqrt{12}\times\sqrt{12}$ phase and the $\mathbb{Z}_{2}$ QSL phase in the ground state phase diagram of the model, and Sec. III.3, on the other hand, focuses on the EE results from our DMRG and for the first time in literature, we show the topological EE inside the $\mathbb{Z}_{2}$ QSL away from the exactly solvable RK point. Finally, Sec. IV reiterates the reason for the success in our symmetrization of the local constraints and points out other interesting constrained quantum lattice models that are either of fundamental theoretical importance or of immediate experimental relevance in various quantum simulators that are waiting to be explored with our algorithm.
@@
Figure 1: QDM on a plaquette as an algorithmic demonstration. (a) The system with black (red) numbers indexing the bonds (sites). (b) The Hamiltonian in terms of local tensors. (c) Schematic of the quantum numbers flows for the MPO in (b). The four numbers are the quantum numbers (number of dimers on each site), and each four numbers set label a virtual bond space. (d) A $6\times 6$ triangular lattice with OBC (PBC) along $x$- ($y$-) direction. The local degree of freedom lives on every bond, and they are labeled according to the DMRG path. For a two-site update on bonds 51 & 52 (green), only 8 quantum numbers (sites in red) are active. All bonds before 51 are colored in blue.
@@ II Method
where the $t$-term flips parallel dimers to other directions in the plaquette and the positive (negative) $V$-term repulses (attracts) parallel dimers. The RK point located at $t=V$ is an exact-solvable point, where the ground state wavefunction consists of the equal superposition of all possible dimer configurations under the constraint [37, 8].
Note that every term in the Hamiltonian commutes with the operator counting the number of dimers connected to each site. Although each kinetic ($t$) term swaps dimer on bonds with different orientations, the four sites it touches are connected to the same numbers of dimers before and after. Consequently, the number of dimer connected to each site is a conserved quantity, i.e., the local constraint, and we have $N_{\textrm{site}}$ many of them defined on the lattice. One can then implement these conserved quantities in the DMRG computation, the same as how one considers the total $S^{z}$ in the Heisenberg model as a good quantum number. Once these symmetries are implemented, one can speed up all computation by making use of the sparse block structure of local tensors in MPO and MPS, and the constraint will be set by the initial MPS and unchanged throughout the computation.
Moreover, the Hamiltonian commutes with operator $M=\prod_{i\in l_{M}}(-1)^{\hat{n}_{i}}$ leading to two topological sectors, where $l_{M}$ is one horizontal path cutting through the whole system (red dashed line in Fig. 3 (b) & (c)). The commutation relation follows from the fact that every plaquette operator in the Hamiltonian touches either 0 or 2 bonds on $l_{M}$, leaving $M$ unchanged. Since we only have PBC along the $y$-direction, this geometry holds only two topological sectors characterized by $M=\pm 1$. This global symmetry can also be implemented in DMRG by tracking a quantity that changes sign upon crossing a dimer located on bonds along $l_{M}$.
@@
$H=J\sum_{\langle i,j\rangle}\bm{S}_{i}\cdot\bm{S}_{j}=J\sum_{\langle i,j\rangle}\{S^{z}_{i}S^{z}_{j}+\frac{1}{2}(S^{+}_{i}S^{-}_{j}+S^{-}_{i}S^{+}_{j})\}$ (2)
@@
In an MPO, the charge flow is marked when constructing the Hamiltonian MPO in the automata process [61, 62, 63, 64, 65, 66]. Suppose one would like to put in the interaction $S^{+}_{2}S^{-}_{3}$, where 2 and 3 are two nearest neighbors in Eq. (2). In the second local tensor in the MPO train, an $S^{+}_{2}$ is put in a block labeled 0 and +1 for charges flowing in and out in the virtual bond spaces, respectively, and $S^{-}_{3}$ is put in the third local tensor, with labels +1 and 0. Since all the terms in the Hamiltonian preserve such symmetry, every term will end with charge 0. By labeling flows for all dimensions like this, all local tensors are block diagonalized, where nonzero elements only appear in blocks where charges flowing in and out are the same, thus, SVD can be performed on each block successively. For the MPS, the charge flow in each local tensor is marked similarly, and the charge flowing out from the last tensor is the total $S^{z}$ for the state.
@@
For QDM or QLM, as mentioned before, the Hamiltonian preserves the number of dimers connected to each site. Therefore, we can label the MPO and MPS similarly to how it is done in the Heisenberg model, with each virtual bond space being labeled with $N_{\mathrm{site}}$ quantum numbers. For the last tensor of MPO, the quantum numbers will flow to $(0,0,\cdots,0)$, similar to the Heisenberg case. For the MPSs, the last flowing out should represent the number of dimers connected to each site in the state, which should be $(1,1,\cdots,1)$ for QDM and $(2,2,\cdots,2)$ for QLM, respectively.
@@
where Pauli matrices $\sigma_{i}^{+}$, $\sigma^{-}_{i}$ acting on the two-dimensional local Hilbert space on the $i$-th bond creates and annihilates a dimer respectively, and $\hat{n}_{i}$ counts the number of dimers on the $i$-th bond. The MPO train formulation for this Hamiltonian is shown in Fig. 1 (b). One can recover the Hamiltonian by contracting all virtual bond dimensions.
@@
Consider the first term in the Hamiltonian $-t\sigma_{1}^{+}\sigma_{2}^{+}\sigma_{3}^{-}\sigma_{4}^{-}$, which is carried in the second channels (columns) in local tensors $M_{1}$,$M_{2}$ and $M_{3}$. $-t\sigma_{1}^{+}$ acting on the first bond creates a dimer here, thus increase the quantum numbers (numbers of dimer) on sites 1 and 2 (Fig. 1 (a)) by one, thus leads to quantum numbers flowing out at the second channel being (1,1,0,0). In $M_{2}$, $\sigma_{2}^{+}$ is placed at the (2,2) block, which means the quantum numbers flowing in is (1,1,0,0) (the quantum numbers flowing out of the second channel in $M_{1}$). A creation on bond 2 will increase quantum numbers on sites 3 and 4 by 1, thus leading to quantum numbers flowing out at the second channel being (1,1,1,1). One can follow any of the terms and check that all quantum numbers indeed flow to (0,0,0,0) eventually (Fig. 1 (c)), as guaranteed by the symmetries of Hamiltonian. More details on the implementation can be found in Appendix A.
In a larger system with more sites or a more complicated lattice structure, the same symmetry can be implemented. However, in practice, one does not need to keep track of all $N_{\textrm{site}}$ many quantum numbers for the whole process, but only the active ones, which are the ones touched by previous bonds at least once and will be touched later. Consider a $6\times 6$ triangular lattice, as shown in Fig. 1 (d), where the numbers indicate the sequence of DMRG sweep, and blue (black) bonds represent the visited (unvisited) bonds. If a two-site update is performed on local tensors 51 & 52 (bonds in green), only 8 quantum numbers (sites in red) are active. For sites connected to only blue bonds, their corresponding quantum numbers cannot be modified anymore and are equal to the values set by the constraint, which are the same throughout all virtual bond dimensions. For sites connected to only black bonds, they have yet to be touched by any dimer, and their corresponding quantum numbers are all 0. Only those sites on the boundary of blue and black bonds carry active quantum numbers. For DMRG simulation on a 2D lattice wrapped around a cylinder, the number of such non-trivial quantum numbers to be stored at each local tensor can be reduced to roughly the number of sites along the periodic boundary condition ($L_{y}$).
@@ III Results
It is expected that at negative $V$, the dimers will form a columnar phase with translational symmetry breaking in one of the three lattice directions (related by 120 rotation) and 2-fold degeneracy in each of the symmetry-breaking directions. The formation of the columnar phase is due to the interplay of the attractive $V$ and an order-by-disorder effect [67, 9, 11, 68] (to overcome the extensive degeneracy $\exp(L_{y})$ generated by sliding each row of dimers at the ordered direction freely). As $V$ moves towards positive values, the columnar phase will transit into a valence bond solid (VBS) with $\sqrt{12}\times\sqrt{12}$ unit cell [11]. Such an enlarged unit cell contains 3 resonating rhombi, i.e., each unit cell contains 12 lattice sites, and it is supposed to have a Bragg peak of the dimer-dimer correlation at the $\bm{X}$ point of the Brillouin zone (BZ). The transition between columnar and VBS phase is suggested to be first order [11, 21].
@@
Figure 4: Dimer correlation of QDM on triangular lattice. (a) Dimer correlation function $S(\bm{k})$ along the high symmetry path (denoted as the black path in panels (b-d)) for $V=-1$, $V=-0.2$ and $V=0.9$. The system size is $L_{x}=24$ and $L_{y}=6$. (b-d) are the correlation in the entire BZ. Inside the columnar phase with $V=-1$, the Bragg peak of the order is at the $\bm{M}$ points (red circle); inside the $\sqrt{12}\times\sqrt{12}$ VBS phase with $V=-0.2$, the order wave vector is at $\bm{X}$ points (red corss), and the $\bm{M}$ point becomes secondary; inside the $\mathbb{Z}_{2}$ QSL phase with $V=0.9$, the dimer correlation is flat as the dimer (as well as vison) correlation are all short-ranged in real-space [24]. Panels (b-d) share the same colorbar on the right.